OpenStudy (anonymous):

A triangle has a base of 13 meters and a height of 6 meters. What is the area of the triangle? A. 78 square meters B. 169 square meters C. 156 square meters D. 39 square meters

OpenStudy (igreen):

A = 1/2bh A = 1/2(13)(6) Multiply

OpenStudy (anonymous):

39

OpenStudy (igreen):

Correct

OpenStudy (anonymous):

o.o So simple I guess

OpenStudy (anonymous):

@Baby_Bear69 K, let's try this again xD

OpenStudy (cggurumanjunath):



OpenStudy (baby_bear69):

alright! haha